# Negative Yields Positive: Unified Dual-Path Adapter for Vision-Language Models

### v1

@@ 4 Experiments
In this section, we conduct extensive experiments on two tasks across 15 datasets. These results demonstrate that our DualAdapter outperforms other state-of-the-art methods, exhibiting enhanced adaptation and generalization capabilities.
@@ 4.1 Experimental Settings
Tasks and Datasets. To validate the effectiveness of our proposed DualAdapter, we evaluate our proposed method on two standard benchmarking tasks: few-shot learning and domain generalization, respectively. For few-shot learning task, we comprehensively evaluate our method on 11 well-known image classification benchmarks: ImageNet [4], Caltech101 [11], OxfordPets [45], StandfordCars [29], Flowers102 [44], Food101 [1], FGVCAircraft [39], DTD [3], SUN397 [71], EuroSAT [18], and UCF101 [57]. For domain generalization, we investigate the generalization capability of our DualAdapter on 4 variant datasets of ImageNet: ImageNet-V2 [50], ImageNet-Sketch [65], ImageNet-A [20], and ImageNet-R [19].
Implementation Details. Following previous works [83, 79], we adopt ResNet-50 [17] backbone as the visual encoder of CLIP in our experiments by default. We adopt prompt ensembling, leveraging textual prompts from both CLIP [48] and CuPL [46] to enhance model performance. For the negative prompts we used for each dataset, please kindly refer to the Supplementary Materials. We set the hyper-parameters $\lambda$ and $\tau$ as $0.75$ and $1$, respectively. Our DualAdapter is trained using the AdamW [36] optimizer with a cosine scheduler [35]. The batch size is set to 256. For $\mathcal{R}_{\mathsf{T}}^{+}$ and $\mathcal{R}_{\mathsf{V}}^{+}$, the learning rate is set to $0.0001$, while for $\mathcal{R}_{\mathsf{T}}^{-}$ and $\mathcal{R}_{\mathsf{V}}^{-}$, the learning rate is set to $0.0005$. Additionally, our model is trained for 200 epochs on the EuroSAT [18] dataset, and for 20 epochs on all other datasets. To ensure the reliability of our results, we perform each experiment three times using different initialization seeds and report the mean accuracy achieved. All experiments are conducted on a single NVIDIA RTX 6000 Ada GPU.
@@
4.2 Few-Shot Learning
In Fig. 5, we compare the few-shot learning performance of our DualAdapter with other state-of-the-art methods on 11 image classification datasets. In the top-left sub-figure, we also present the average classification accuracy across all 11 datasets. The results indicate that our DualAdapter consistently outperforms other methods across various few-shot learning settings by substantial margins. Moreover, our DualAdapter demonstrates more pronounced performance improvements in specialized classification tasks, such as satellite image and texture classification on the EuroSAT [18] and DTD [3] datasets. With 16-shot training on these two datasets, our DualAdapter surpasses Tip-Adapter-F [79] by a notable 3.56% and 3.50%, and enhances the performance of zero-shot CLIP [48] significantly by 45.56% and 30.50%, respectively. For full numerical results, please refer to our Supplementary Materials. Overall, the consistently superior performance on 11 datasets fully demonstrates the effectiveness of our DualAdapter.
Figure 5: Performance comparisons on few-shot learning on 11 image classification datasets. For each dataset, we report the mean accuracy and 95% confidence interval over 3 random seeds of our DualAdapter on 1-/2-/4-/8-/16-shot settings.
4.3 Robustness to Natural Distribution Shifts
@@
DualAdapter (Ours) 66.52 57.87 36.38 25.73 61.12 45.28
In Table 1, we compare the generalization performance of our DualAdapter with other state-of-the-art methods in the presence of distribution shifts. Specifically, all the models are trained solely on 16-shot ImageNet [4], and directly tested on 4 out-of-distribution ImageNet variant datasets. As shown in Table 1, our DualAdapter not only achieves state-of-the-art performance on the source dataset but also attains an average performance gain of 1.18% across 4 out-of-distribution (OOD) target datasets. These experimental results demonstrate that by enabling adaptation through both positive and negative paths, our DualAdapter can more effectively prevent overfitting to the source dataset, thereby enhancing robustness against distribution shifts.
4.4 Ablation Studies
[…]
Table 2: Ablation studies for different variants of DualAdapter. We evaluate the few-shot adaptation capabilities of four DualAdapter variants on ImageNet [4].
# Method $\mathcal{R}_{\mathsf{T}}^{+}$ $\mathcal{R}_{\mathsf{T}}^{-}$ $\mathcal{R}_{\mathsf{V}}^{+}$ $\mathcal{R}_{\mathsf{V}}^{-}$ 1-shot 2-shot 4-shot 8-shot 16-shot
[…]
3 ${}^{+}$ 62.83 63.31 63.95 65.13 66.27
4 ${}^{-}$ 62.65 63.07 63.60 64.36 65.12
5 DualAdapter 62.89 63.47 64.12 65.37 66.52
Effects of Different Visual Backbones. We also implement our DualAdapter with various visual encoders, including ResNet [17] and ViT [6], and evaluate its performance against other fine-tuning approaches in Fig. 6 (Left). We can see that our DualAdapter consistently exceeds other methods across all visual backbones, indicating the general effectiveness of our approach.
Figure 6: More ablation results. (Left) Performance comparison of our DualAdapter with other methods on few-shot learning task using different visual backbones; (Middle) Sensitivity analysis of $\lambda$ from Eq. (10) on ImageNet [4]; (Right) Sensitivity analysis of $\tau$ from Eq. (12) on 3 datasets, with the value of $\tau$ presented on a logarithmic scale.
[…]
Scalability to More Shots. When scaling up to more than 16 shots, Tip-Adapter-F [79] limits the cache size to prevent memory overflow. In contrast, our DualAdapter features reduced memory requirements, enabling it to operate without imposing cache size constraints, even in the 128-shot setting. Performance comparisons in Table 4.4 indicate that our DualAdapter outperforms Tip-Adapter-F [79] more significantly in higher shot settings, e.g., by 1.24% in the 128-shot setting.
Efficiency Comparison. We also compare the efficiency of DualAdapter with existing methods in Table 4.4. Our DualAdapter achieves the highest accuracy while also exhibiting advantageous computational efficiency: (1) Compared to prompt-based learning methods such as CoOp [84] and ProGrad [87], our DualAdapter requires approximately 300$\times$ less training time and over 1000$\times$ fewer FLOPs since we do not need to propagate gradients through the textual encoder; (2) Compared to adapter-style fine-tuning methods, our DualAdapter is also competitive in efficiency. It requires 10$\times$ less training time than CLIP-Adapter [12], and demands 3$\times$ fewer FLOPs and 4$\times$ fewer learnable parameters than Tip-Adapter-F [79].
Table 3: Performance comparison with Tip-Adapter(-F) [79] on ImageNet [4] with 16-/32-/64-/128-shot settings.
Table 4: Efficiency comparison with other existing methods on 16-shot ImageNet [4].
Method 16-shot 32-shot 64-shot 128-shot
CLIP [48] ${}^{\dagger}$0-shot: 60.33
Tip-Adapter [79] 62.03 62.51 62.88 63.15
Tip-Adapter-F [79] 65.47 66.58 67.96 69.74
DualAdapter (Ours) 66.52 67.68 69.01 70.98
@@ 4.4 Ablation Studies
DualAdapter (Ours) 5 min 20 0.009 4.10M 66.52
Table 4: Efficiency comparison with other existing methods on 16-shot ImageNet [4].
@@ 5 Conclusion
In this work, we introduce DualAdapter, a novel approach to effectively adapt vision-language models for downstream tasks by integrating both positive and negative adapters. This innovative approach allows for effective few-shot adaptation from a dual perspective. Furthermore, we develop an unsupervised similarity-based label refinement approach to mitigate the adverse effects of low-quality image samples during the few-shot adaptation process. Comprehensive evaluations on 15 diverse datasets demonstrate that DualAdapter outperforms the state-of-the-art methods in both few-shot learning and domain generalization tasks.
Limitations. We identify two potential limitations of our DualAdapter: (1) Its efficacy in zero-shot situations is constrained due to the scarcity of original negative descriptions in the CLIP training corpus; (2) Similar to other adapter-style approaches such as Tip-Adapter-F [79], our DualAdapter fine-tuned on a specific task cannot be directly applied to another task without additional adaptation. However, recently, Wang et al. [68] shows that adapter-style fine-tuning methods can be extended for these scenarios using the $k$NN algorithm.
Acknowledgement
This work has been funded in part by the Army Research Laboratory (ARL) under grant W911NF-23-2-0007 and W911NF-19-2-0146, and the Air Force Office of Scientific Research (AFOSR) under grants FA9550-18-1-0097 and FA9550-18-1-0251.
@@ Acknowledgement
In this supplementary document, we provide additional details and experimental results to enhance understanding and insights into our proposed DualAdapter. This supplementary document is organized as follows:
@@
    We compare our DualAdapter with other state-of-the-art methods in domain generalization tasks, utilizing an alternative ViT backbone, in Section A.2.
@@
    Detailed statistics for all utilized datasets are provided in Section B.1.
@@
    We list the specific positive and negative prompts we used for each dataset in Section B.2.
@@
    Finally, we explore potential avenues for future research in Section C.
@@ A.1 Full Numerical Results on Few-Shot Learning
In Fig. 5 in Section 4.2 of the main text, we have evaluated our DualAdapter on few-shot learning task and compared with other state-of-the-art methods. In Table A1, we present the corresponding full numerical results on the few-shot learning task. We also report the 95% confidence interval over 3 random seeds of our DualAdapter to ensure reliability of our results. In the last column, we present the average recognition accuracy over 11 datasets. The results indicate that our DualAdapter consistently outperforms other state-of-the-art methods across various few-shot learning settings by substantial margins.
Our DualAdapter demonstrates superior recognition performance across nearly all tested scenarios, with certain exceptions in lower-shot settings for the Food101 [1] and OxfordPets [45] datasets. We attribute this to the prevalent challenge of overfitting, a common issue not exclusive to our approach but also affecting many existing methods, especially TaskRes [75]: While TaskRes consistently secures the second-best performance across the other 9 datasets, it underperforms significantly on these two. We hypothesize that this issue arises from the noisy training data with intense colors and sometimes wrong labels [1, 45]. However, in this work, we have designed a label refinement mechanism specifically to address this issue. As a result, our DualAdapter secures a relatively robust performance and achieves the second-best on these two datasets.
Table A1: Full numerical results on few-shot learning task. For each dataset, we report the mean accuracy and 95% confidence interval over 3 random seeds of our DualAdapter on 1-/2-/4-/8-/16-shot settings. ${}^{\dagger}$e report the zero-shot performance of CLIP [48] for all settings. For TaskRes [75], we report the results using the enhanced base classifier (i.e., TaskRes*). The best results are in bold and the second are underlined.
@@
${}^{\dagger}$ero-shot CLIP [48] 1-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
@@
DualAdapter (Ours) (±$\pm 0.33$) (±$\pm 0.48$) (±$\pm 0.47$) (±$\pm 0.18$) (±$\pm 0.34$) (±$\pm 0.22$) (±$\pm 0.13$) (±$\pm 0.31$) (±$\pm 0.28$) (±$\pm 0.08$) (±$\pm 0.23$) (±$\pm 0.28$)
${}^{\dagger}$ero-shot CLIP [48] 2-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
@@
DualAdapter (Ours) (±$\pm 0.26$) (±$\pm 0.31$) (±$\pm 0.87$) (±$\pm 0.37$) (±$\pm 0.18$) (±$\pm 0.17$) (±$\pm 0.06$) (±$\pm 0.28$) (±$\pm 0.21$) (±$\pm 0.33$) (±$\pm 0.48$) (±$\pm 0.31$)
${}^{\dagger}$ero-shot CLIP [48] 4-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
@@
DualAdapter (Ours) (±$\pm 0.21$) (±$\pm 0.49$) (±$\pm 0.59$) (±$\pm 0.29$) (±$\pm 0.25$) (±$\pm 0.11$) (±$\pm 0.17$) (±$\pm 0.26$) (±$\pm 0.36$) (±$\pm 0.29$) (±$\pm 0.51$) (±$\pm 0.32$)
${}^{\dagger}$ero-shot CLIP [48] 8-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
@@
DualAdapter (Ours) (±$\pm 0.23$) (±$\pm 0.55$) (±$\pm 0.54$) (±$\pm 0.63$) (±$\pm 0.26$) (±$\pm 0.21$) (±$\pm 0.09$) (±$\pm 0.21$) (±$\pm 0.51$) (±$\pm 0.15$) (±$\pm 0.39$) (±$\pm 0.34$)
${}^{\dagger}$ero-shot CLIP [48] 16-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77

### v2

@@ 4 Experiments
In this section, we conduct extensive experiments on two tasks across 15 datasets. These results demonstrate that our proposed method is simple yet highly effective, surpassing other state-of-the-art methods in both few-shot adaptation and domain generalization capabilities.
@@ 4.1 Experimental Settings
Figure 5: Performance comparisons on few-shot learning on 11 image classification datasets. For each dataset, we report the mean accuracy and 95% confidence interval over 3 random seeds of our SimNL on 1-/2-/4-/8-/16-shot settings.
Tasks and Datasets. To validate the effectiveness of SimNL, we evaluate our method on two standard benchmarking tasks: few-shot learning and domain generalization, respectively. For few-shot learning task, we comprehensively evaluate our method on 11 well-known image classification benchmarks: ImageNet [6], Caltech101 [9], OxfordPets [35], StandfordCars [24], Flowers102 [33], Food101 [2], FGVCAircraft [30], DTD [5], SUN397 [63], EuroSAT [14], and UCF101 [48]. For domain generalization, we evaluate the generalizability of our SimNL on 4 variants of ImageNet: ImageNet-V2 [40], ImageNet-Sketch [55], ImageNet-A [16], and ImageNet-R [15]. Moreover, we also explore the adaptation of VLMs using a noisy few-shot sample set from ImageNet [6], where labels are randomly flipped to simulate real-world scenarios.
Implementation Details. Following previous works, we adopt ResNet-50 [13] backbone as the visual encoder of CLIP in our experiments by default. We adopt prompt ensembling, leveraging textual prompts from CLIP [38] to enhance model performance. For the negative prompts we used for each dataset, please kindly refer to Appendix C.2. We set the hyper-parameters $\lambda$ and $\tau$ as $0.75$ and $1$, respectively. Our SimNL is trained using the AdamW [29] optimizer with a cosine scheduler [28]. The batch size is set to 256. For $\mathcal{R}_{\mathsf{T}}^{+}$ and $\mathcal{R}_{\mathsf{V}}^{+}$, the learning rate is set to $0.0001$, while for $\mathcal{R}_{\mathsf{T}}^{-}$ and $\mathcal{R}_{\mathsf{V}}^{-}$, the learning rate is set to $0.0005$. Our model is trained for 200 epochs on the EuroSAT [14] dataset, and for 20 epochs on all other datasets. To ensure the reliability of our results, we perform each experiment three times using different initialization seeds and report the mean accuracy achieved. All experiments are conducted on a single 48GB NVIDIA RTX 6000 Ada GPU.
@@
4.2 Results and Analysis
Few-Shot Learning. In Figure 5, we compare the few-shot learning performance of our proposed method with other state-of-the-art methods on 11 image classification datasets. In the top-left sub-figure, we also present the average classification accuracy across all 11 datasets. The results indicate that our method consistently outperforms other methods across various few-shot learning settings by substantial margins. Moreover, our proposed method demonstrates more pronounced performance improvements in specialized classification tasks, such as satellite image and texture classification on the EuroSAT [14] and DTD [5] datasets. With 16-shot training on these two datasets, our method surpasses Tip-Adapter-F [71] by a notable 3.56% and 3.50%, respectively. For full numerical results, please refer to Table A1 in Appendix A.1. Overall, the consistently superior performance on 11 datasets fully demonstrates the general effectiveness of our proposed approach.
Robustness to Natural Distribution Shifts. In Table 1, we compare the generalizability of our SimNL with other methods in the presence of distribution shifts. Specifically, all the models are trained solely on 16-shot ImageNet [6], and directly tested on 4 out-of-distribution ImageNet variant datasets. As shown in Table 1, our method not only achieves state-of-the-art performance on the source dataset but also attains an average performance gain of 1.18% across 4 out-of-distribution (OOD) target datasets. These experimental results indicate that by enabling adaptation from both positive and negative perspectives, our method enhances robustness against distribution shifts.
@@ 4.2 Results and Analysis
SimNL (Ours) 66.52 57.87 36.38 25.73 61.12 45.28
Robustness to Label Noise. In Table 2, we report the performance of Tip-Adapter-F [71] and our SimNL on a noisy 16-shot ImageNet [6], where we randomly flip 10% to 50% labels. As shown, applying our instance reweighting enhances both Tip-Adapter-F and SimNL’s robustness to label noise. Notably, our reweighting technique also improves performance when no label noise is introduced, as it can identify and downweight outliers during adaptation.
Table 2: Comparison of robustness to label noise on noisy 16-shot ImageNet [6]. We apply our instance reweighting technique to both Tip-Adapter-F [71] and our SimNL, and report the performance across four levels of noise.
Method 0% 10% 25% 50%
Tip-Adapter-F [71] 65.52 64.93 64.04 62.47
+ Reweighting 65.64 65.25 64.55 63.39
Performance Gain +0.12 +0.32 +0.51 +0.92
SimNL (Ours) 66.31 65.54 64.77 63.37
+ Reweighting 66.52 65.82 65.16 64.02
Performance Gain +0.21 +0.28 +0.39 +0.65
Figure 6: More ablation results. (Left) Performance comparison of our SimNL with others on few-shot learning using different visual backbones; (Middle) Sensitivity analysis of $\lambda$ from Eq. (9) on ImageNet [6]; (Right) Sensitivity analysis of $\tau$ from Eq. (11) on 3 datasets.
Efficiency Comparison. We also compare the efficiency of SimNL with existing methods in Table 3. Our method achieves the highest accuracy while also exhibiting advantageous computational efficiency: (1) Compared to prompt-based learning methods such as CoOp [74] and ProGrad [75], our proposed method requires approximately 300$\times$ less training time and over 1000$\times$ fewer FLOPs since we do not need to propagate gradients through the textual encoder; (2) Compared to adapter-style fine-tuning methods, our SimNL requires 10$\times$ less training time than CLIP-Adapter [10], and demands 3$\times$ fewer FLOPs than Tip-Adapter-F [71] and GraphAdapter [25].
Table 3: Efficiency comparison with other existing methods on 16-shot ImageNet [6]. We report the training time, number of epochs, FLOPs, and the number of parameters for each method.
@@
GraphAdapter [25] 5 min 20 0.030 4.15M 65.70
SimNL (Ours) 5 min 20 0.009 4.15M 66.52
4.3 Ablation Studies
Effectiveness of Different Components. In Table 4, we conduct a systematic analysis of the impacts of various components within our SimNL framework. More specifically, we assess the performance of four distinct SimNL variants, each configured to allow two learnable residues to be updated while keeping the others fixed. We observe that the textual variant (SimNL-T) and the positive variant (SimNL-P) generally exhibit superior efficiency compared to their visual and negative counterparts. The full SimNL method, which integrates all four branches, surpasses the individual variants by achieving the highest average accuracy of 77.80% in the 16-shot scenario.
Table 4: Ablation studies for different variants of our method. We evaluate the few-shot adaptation capabilities of four variants and report their average performance across all 11 datasets.
Method $\mathcal{R}_{\mathsf{T}}^{+}$ $\mathcal{R}_{\mathsf{T}}^{-}$ $\mathcal{R}_{\mathsf{V}}^{+}$ $\mathcal{R}_{\mathsf{V}}^{-}$ 1-shot 2-shot 4-shot 8-shot 16-shot
SimNL-T 67.08 69.70 71.86 74.17 77.08
SimNL-V 64.60 66.03 68.25 71.84 74.41
SimNL-P 66.72 69.30 71.49 73.98 76.76
SimNL-N 66.23 68.10 70.55 72.33 74.17
SimNL 67.45 70.06 72.43 75.06 77.80
Effects of Different Visual Backbones. We also implement our SimNL with various visual encoders, including ResNet [13] and ViT [8], and evaluate their performance against other adaptation methods in Figure 6 (Left). We can see that our SimNL consistently exceeds other methods across all visual backbones, indicating the general effectiveness of our negative learning approach.
Sensitivity Analysis of Hyper-Parameters. We provide sensitivity analysis for the hyper-parameters $\lambda$ and $\tau$ in Figure 6. The hyper-parameter $\lambda$ from Eq. (9) controls the combination of the positive and negative predictions. In Figure 6 (Middle), we can observe that setting $\lambda=0.75$ consistently yields the optimal performance. Moreover, in Figure 6 (Right), we adjust the temperature hyper-parameter $\tau$ in Eq. (11) from $0.04$ to $25$ and report the 16-shot accuracy of our method. We can see that by reweighting the few-shot samples, we achieve performance improvements ranging from 0.21% to 0.55% across three datasets.
Scalability to More Shots. In Table 5, we demonstrate that our SimNL scales effectively to more-shot settings, consistently outperforming Tip-Adapter-F [71] by 1.05% to 1.24% as the number of shots increases.
Table 5: Scalability to more shots. We show performance comparison with Tip-Adapter(-F) on ImageNet at higher-shot settings.
Method 16-shot 32-shot 64-shot 128-shot
CLIP [38] 0-shot: 60.33
Tip-Adapter [71] 62.03 62.51 62.88 63.15
Tip-Adapter-F [71] 65.47 66.58 67.96 69.74
SimNL (Ours) 66.52 67.68 69.01 70.98
@@ 5 Conclusion
In this work, we introduce SimNL, a simple and effective approach that applies the concept of negative learning to vision-language few-shot adaptation. Specifically, we transform the open-vocabulary CLIP model into a negative classifier, which is further optimized to exclude incorrect classes based on the image input. During inference, we conduct simultaneous positive classification and negative exclusion to enhance the overall prediction accuracy. Furthermore, we develop an unsupervised few-shot instance reweighting approach to mitigate the adverse effects of noisy image samples during few-shot adaptation . Comprehensive evaluations on 15 diverse datasets demonstrate that our proposed SimNL outperforms the state-of-the-art methods in both few-shot learning and domain generalization tasks, while maintaining competitive efficiency.
Acknowledgements. This work has been funded in part by the Army Research Laboratory (ARL) award W911NF-23-2-0007, DARPA award FA8750-23-2-1015, and ONR award N00014-23-1-2840.
@@
In appendix, we provide additional details and experimental results to enhance understanding and insights into our proposed SimNL. This supplementary document is organized as follows:
@@
    We compare our SimNL with other state-of-the-art methods in domain generalization tasks, utilizing an alternative ViT backbone, in Section A.2.
@@
    We differentiate our method from related work in Section B.
@@
    Detailed statistics for all utilized datasets are provided in Section C.1.
@@
    We present the specific positive and negative prompts we used for each dataset in Section C.2.
  • $\bullet$
    We list the license information for all used assets in Section D.
  • $\bullet$
    Finally, we explore potential future work and discuss the limitations and broader impacts of this work in Section E.
@@ A.1 Full Numerical Results on Few-Shot Learning
In Figure 5 of the main text, we have evaluated our SimNL on few-shot learning task and compared with other state-of-the-art methods. In Table A1, we present the corresponding full numerical results on the few-shot learning task. We also report the 95% confidence interval over 3 random seeds of our SimNL to ensure reliability of our results. In the last column, we present the average recognition accuracy over 11 datasets. The results indicate that our SimNL consistently outperforms other state-of-the-art methods across various few-shot learning settings by substantial margins.
Our SimNL demonstrates superior recognition performance across nearly all tested scenarios, with certain exceptions in lower-shot settings for the Food101 [2] and OxfordPets [35] datasets. We attribute this to the prevalent challenge of overfitting, a common issue not exclusive to our approach but also affecting many existing methods, especially TaskRes [65]: While TaskRes consistently secures the second-best performance across the other 9 datasets, it underperforms significantly on these two. We hypothesize that this issue arises from the noisy training data with intense colors and sometimes wrong labels [2, 35]. However, in this work, we have designed a label refinement mechanism specifically to address this issue. As a result, our SimNL secures a relatively robust performance and achieves the second-best on these two datasets.
Table A1: Full numerical results on few-shot learning task. For each dataset, we report the mean accuracy and 95% confidence interval over 3 random seeds of our SimNL on 1-/2-/4-/8-/16-shot settings. We report the zero-shot performance of CLIP [38] for all settings. For TaskRes [65], we report the results using the enhanced base classifier (i.e., TaskRes*). The best results are in bold and the second are underlined.
@@
Zero-shot CLIP [38] 1-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
@@
SimNL (Ours) (±$\pm 0.33$) (±$\pm 0.48$) (±$\pm 0.47$) (±$\pm 0.18$) (±$\pm 0.34$) (±$\pm 0.22$) (±$\pm 0.13$) (±$\pm 0.31$) (±$\pm 0.28$) (±$\pm 0.08$) (±$\pm 0.23$) (±$\pm 0.28$)
Zero-shot CLIP [38] 2-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
@@
SimNL (Ours) (±$\pm 0.26$) (±$\pm 0.31$) (±$\pm 0.87$) (±$\pm 0.37$) (±$\pm 0.18$) (±$\pm 0.17$) (±$\pm 0.06$) (±$\pm 0.28$) (±$\pm 0.21$) (±$\pm 0.33$) (±$\pm 0.48$) (±$\pm 0.31$)
Zero-shot CLIP [38] 4-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
@@
SimNL (Ours) (±$\pm 0.21$) (±$\pm 0.49$) (±$\pm 0.59$) (±$\pm 0.29$) (±$\pm 0.25$) (±$\pm 0.11$) (±$\pm 0.17$) (±$\pm 0.26$) (±$\pm 0.36$) (±$\pm 0.29$) (±$\pm 0.51$) (±$\pm 0.32$)
Zero-shot CLIP [38] 8-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
@@
SimNL (Ours) (±$\pm 0.23$) (±$\pm 0.55$) (±$\pm 0.54$) (±$\pm 0.63$) (±$\pm 0.26$) (±$\pm 0.21$) (±$\pm 0.09$) (±$\pm 0.21$) (±$\pm 0.51$) (±$\pm 0.15$) (±$\pm 0.39$) (±$\pm 0.34$)
Zero-shot CLIP [38] 16-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77